# Probing Non-Fermi-Liquid Behaviour of Composite Fermi Liquid via Efficient Thermal Simulations

### v1

@@ Abstract
The two-dimensional electron gas in a perpendicular magnetic field, i.e., the quantum Hall system, is remarkably rich. At half filling of the lowest Landau level, it has been predicted that “composite fermions”—emergent quasiparticle of an electron with two magnetic flux quanta—can experience zero net magnetic field and form a Fermi sea, dubbed composite Fermi liquid (CFL). However, the seemingly simple appearance of CFL is a strongly correlated quantum many-body state in disguise, and to solve it in a controlled manner is extremely difficult, to the level that the thermodynamic properties of CFL is still largely unknown. In this work, we perform state-of-the-art thermal tensor network simulations on the $\nu=1/2$ Landau level systems, and observe low-temperature power-law behaviour of the specific heat, signaling the gapless nature of CFL. More importantly, the power is extracted to be closed to $2/3$, clearly deviated from the ordinary linear-$T$ Fermi liquid behaviour, suggesting the coupling between the CFs and the dynamical emergent gauge field and therefore revealed the quantum many-body aspect of the CFL state. Relevance of our methodology to other quantum Hall settings and moiré systems is discussed.
@@
Introduction.— The two-dimensional electron gas subjected to a perpendicular magnetic field—known as the quantum Hall (QH) system—exhibits remarkable physical richness. In the integer quantum Hall regime, the Hall conductance $\sigma_{xy}$ is quantized precisely in integer multiples of $e^{2}/h$ [1]. In the fractional quantum Hall regime, $\sigma_{xy}$ quantizes at fractional values of $e^{2}/h$, accompanied by the emergence of fractionally charged quasiparticles obeying abelian fractional statistics [2, 3, 4]. Notably, at the filling factor $\nu=1/2$, the system exhibits metallic behavior, characterized by a finite longitudinal resistivity $\rho_{xx}$ and the absence of $\rho_{xy}$ pleateau  [5].
In recent years, the fractional Chern insulator (FCI) [6, 7, 8, 9, 10, 11]—a lattice analogue of the fractional quantum Hall effect—has attracted considerable interest in the context of two-dimensional moiré materials, with extensive theoretical explorations [12, 13, 14, 15, 16, 17, 18, 19, 20]. Earlier experiments have reported the FCIs at the fractional filling of the Hofstadter band [21] in a bilayer graphene heterostructure [22], and the fractional filling of the native Chern band in magic-angle twisted bilayer graphene (TBG) with the finite magnetic field to improve the quantum geometry [23, 24]. More recently, the zero-field FCIs are observed in twisted bilayer MoTe2 (tMoTe2) and rhombohedral graphene/hBN superlattices [25, 26, 27, 28, 29]. Notably, the zero-field metallic state at $\nu=1/2$ of the topological bands have also been theoretically and experimentally explored [30, 31, 26, 29].
It is accepted that, while the integer quantum Hall effect is well understood within a non-interacting electron picture [32, 33], the fractional effect stems dominantly from electron-electron interactions. This leads to the formation of an incompressible quantum liquid hosting fractionally charged quasiparticles, first captured by Laughlin’s wave function at fillings $\nu=1/(2m+1)$ [34]. For other filling fractions, a hierarchical construction was proposed by Haldane [35], Halperin [3], and Laughlin [36], wherein quasiparticles themselves will form new incompressible quantum liquid, analogous to the original electron forming Laughlin state.
@@
Alternatively, the composite fermion (CF) theory postulates that electrons attach an even number of magnetic flux quanta, forming emergent quasiparticles that experience a reduced effective magnetic field [37, 38]. Along the line of this flux-attachment concept, mean-field theories [39, 40, 41, 42] are successful: near $\nu=1/2$, ballistic transport experiments are consistent with composite fermions moving in a weak residual magnetic field [43, 44, 45]; at half-filling, the external magnetic field is entirely canceled by the attached flux, resulting in a Fermi sea of composite fermions—a picture corroborated by surface acoustic wave measurements [46]. More recently, this conventional CF picture has been revisited in light of connections to 3D topological insulators (TIs), reformulating composite fermions as massless Dirac fermions reminiscent of TI surface states [47, 48, 49, 50]. This “composite Dirac fermion” theory couples fermions to an emergent gauge field, exhibiting non-Fermi-liquid behavior and providing a field-theoretic realization of (2+1)-dimensional quantum electrodynamics (QED3) as a strongly coupled conformal field theory, with both condensed matter [51, 52, 53, 54, 55, 56, 57, 58] and high-energy [59, 60, 61, 62, 63, 64, 65, 66] physics correspondence.
@@
Numerical studies of the fractional quantum Hall (FQH) effect and composite Fermi liquid (CFL) have employed techniques such as exact diagonalization (ED), variational Monte Carlo (VMC), density matrix renormalization group (DMRG), and time-dependent variational principle (TDVP). In contrast to the well-established spectral and entanglement properties of gapped FCI/FQH phases [67, 68, 69, 70, 71], the numerical characterization of the half-filled CFL remains challenging. Key issues include the role of particle-hole symmetry [72], the scaling of entanglement entropy [73, 74, 75], effects of anisotropy [76], the relationship between the Fermi sea of bare electrons at zero field and that of composite fermions at high field [77], the influence of discrete rotational symmetry $C_{N}$ [78], and the thermoelectric response [79]. In all these directions, new computation and discussions of often-time contradicting results are actively going on.
Despite the intensive investigations, fundamental questions remain: What is the relationship between the hierarchical description and the composite fermion theory? Are they dual representations of the same underlying physics, as argued by Read [80] and Blok and Wen [81], or do they constitute fundamentally distinct frameworks, as suggested by Jain [37]? Is the $\nu=1/2$ state consistent with a Dirac composite Fermi liquid? Is particle-hole symmetry preserved in this state? Particularly, what are the defining thermodynamic signatures of the CFL?
@@
In this work, we investigate the thermodynamic properties of the composite Fermi liquid at half filling of the lowest Landau level using large-scale thermal tensor network simulations, specifically the tangent-space tensor renormalization group (tanTRG) method [82]. We observe a low-temperature specific heat that scales as $T^{\alpha}$, with $\alpha\approx 2/3$—indicative of gapless excitations and markedly inconsistent with the linear-$T$ dependence expected for a conventional Fermi liquid. This power-law behavior provides direct evidence for the strongly correlation nature of CFL state, in that, the coupling between composite fermions and a dynamical emergent gauge field gives rise to non-Fermi liquid properties.
@@
where $V(\mathbf{r}_{1},\mathbf{r}_{2})\equiv\frac{e^{-|\mathbf{r}_{1}-\mathbf{r}_{2}|/\lambda}}{|\mathbf{r}_{1}-\mathbf{r}_{2}|}$ is the Coulomb-Yukawa interaction between electrons, $n(\mathbf{r})\equiv c^{\dagger}(\mathbf{r})c(\mathbf{r})$ is the particle number operator and the factor $\tfrac{1}{2}$ serves to guaranteer the half-filling condition in the grand-canonical ensemble. We then project the above Hamiltonian onto the lowest Landau level (LLL) via $c(\mathbf{r})=\sum_{n}\phi_{n}(\mathbf{r})c_{n}$, where $\phi_{n}(\mathbf{r})\propto e^{-\frac{1}{2}(\frac{x-x_{n}}{\ell})^{2}}e^{\mathrm{i}k_{n}y}$ is the single-particle LLL wavefunction in the cylindrical geometry (i.e., the Landau gauge) with $k_{n}=2\pi n/L_{y}$, $x_{n}=2\pi n\ell^{2}/L_{y}$, $n\in[0,N-1]$ and $N$ the number of LLL orbitals [c.f. Fig. 1 (a)]. We set the magnetic length $\ell=\sqrt{\hbar c/(eB)}=1$ and $k_{\mathrm{B}}=1$ as the unit throughout. This yields the LLL-projected Hamiltonian,
@@
where the form factor $\mathcal{A}_{m,n,k,l}$ involves spatial integration of the interaction $V(\mathbf{r}_{1},\mathbf{r}_{2})$ multiplying with the corresponding LLL wavefunction, which can be more conveniently evaluated by performing Fourier transformation [see Supplementary Materials (SM) [83] for more details].
Tangent-space tensor renormalization group (tanTRG) in the LLL basis.— Here, we briefly recapitulate the idea of tangent-space tensor renormalization group (tanTRG) to investigate the thermodynamic properties for quantum lattice models, and then adapt tanTRG to the LLL orbital basis. Generically, we can construct the matrix product operator (MPO) representation for Hamiltonian $\mathcal{H}$ with local interactions. At a sufficiently high temperature $T=1/\tau$ (small inverse temperature $\tau$), the MPO form of the thermal density operator can then be obtained with high precision by, $\rho(\tau)\equiv e^{-\tau\mathcal{H}}=I-\tau\mathcal{H}$. This can then evolve to lower-temperature regime following the flow equation $\partial\rho/\partial\beta=\mathcal{P}(-\mathcal{H}\rho)\mathcal{P}$ constraint to a fixed-bond-dimension matrix product operator (MPO) manifold $\mathcal{M}_{\mathrm{MPO}}$, with $\mathcal{H}\rho$ being the tangent vector in the full Hilbert space which generally takes the density matrix $\rho$ out of $\mathcal{M}_{\mathrm{MPO}}$, and the projector $\mathcal{P}$ serves to projecting the tangent vector back to the tangent space of the MPO manifold $\mathcal{T}_{\rho}\mathcal{M}_{\mathrm{MPO}}$. For more details concerning the MPO parameterization of the above constraint flow equation, we refer the interested readers to Ref. [82]. This idea is proved to be successful in accessing the challenging low-temperature properties of 2D Hubbard model and $t$-$J$ model for cuprate [82, 84], Bilayer $t$-$J$-$J_{\perp}$ model for Nickelate superconductor [85], and real-space effective model of MoTe2 [86].
@@
Specifically, in this work, we are working in the many-body basis consisted of LLL wavefunctions, $|q_{0},q_{1},\cdots,q_{N-1}\rangle=(c^{\dagger}_{0})^{q_{0}}(c^{\dagger}_{1})^{q_{1}}\cdots(c^{\dagger}_{N-1})^{q_{N-1}}|\Omega\rangle$, with the quantum number $q_{i}$ denoting the number of particles (either 0 or $1$) filled in the $i$th LLL orbital, and $|\Omega\rangle$ being the many-body vacuum. A generic operator $O$ in this many-body LLL basis then express as $O=\sum_{\{q_{i},q_{i}^{\prime}\}}O_{\{q_{i}\},\{q_{i}^{\prime}\}}|\{q_{i}\}\rangle\langle\{q_{i}^{\prime}\}|$, where $O_{\{q_{i}\},\{q_{i}^{\prime}\}}\equiv\langle\{q_{i}\}|O|\{q_{i}^{\prime}\}\rangle$ is a $2^{N}\times 2^{N}$-dimensional tensor, and can be parameterized into the MPO form, $O_{\{q_{i}\},\{q_{i}^{\prime}\}}=\sum_{b_{0},b_{1},\cdots,b_{N}}\prod_{i=0}^{N-1}A_{b_{i},b_{i+1},q_{i},q_{i}^{\prime}}$ with the virtual indices $b_{i}$’s taking values from $1$ to $D$ (bond dimension), as shown in Fig. 1 (b). For the Hamiltonian $\mathcal{H}^{\mathrm{LLL}}$ and thus the initial high-temperature density operator $\rho(\tau)$, we can find an exact MPO form via the finite-size automate machine, and then the lower-temperature density operator $\rho(\beta)$ can then be tracked with the tanTRG protocol.
@@
Benchmark results.— In Fig. 2, we benchmark tanTRG method in a small system size $N=16$. Here, we take the circumference of the cylinder $L_{y}=12$, and the effective screening length $\lambda=10\tfrac{2\pi\ell^{2}}{L_{y}}$. We firstly calculate the free energy per orbital $f(T)\equiv-\frac{1}{N}T\ln Z$ with $Z=\operatorname{Tr}\rho$ the partition function.
In Fig. 2 (a), we have shown three curves of the free energy relative errors $\delta f\equiv|f_{\mathrm{tanTRG}}-f_{\mathrm{ED}}|/|f_{\mathrm{ED}}|$ with incremental bond dimensions $D=400,800,1600$. For all the three curves, at the high-temperature regime, the relative errors remain within the order of $10^{-6}$ as the discarded weights can be neglected and the final error comes from the Trotter error and the projection error (the one results from projecting the full tangent vector into $T_{\rho}\mathcal{M}_{\mathrm{MPO}}$). At intermediate temperature $T\sim 10^{-1}$, the thermal state changes significantly as indicated by the round peak of the specific heat [c.f. the inset of Fig. 2 (b)], and the finite discarded weights set in, resulting in a relatively higher errors of free energies. At lower-temperature regime, finite-size gap sets in and the thermal density operator $\rho$ converges to $|\psi_{0}\rangle\langle\psi_{0}|$ (ground state) which in this example possess lower (purified) entanglement entropy thus lower $\delta f$ relative to the intermediate-$T$ cases. Furthermore, by increasing $D$, the precision increases from $10^{-3}$ to $10^{-5}$ continuously in the low-$T$ regime.
In Fig. 2 (b), we calculate specific heat $c_{V}\equiv\partial e/\partial T$, with $e\equiv\frac{1}{N}\operatorname{Tr}(\mathcal{H}\rho)/Z$ being the internal energy per orbital. Compared with the exact diagonalization results, the relative errors of specific heat $\delta c_{V}$ is shown as function of temperature $T$. Similar to the free energy, the specific heat $c_{V}$ possesses high precision with relative errors within $10^{-8}$ for high temperatures $T\gtrsim 1$, which increase quickly at $\sim 10^{-1}$, and at lower temperatures $\delta c_{V}$ even decreases slightly due to existence of the finite-size gap. Furthermore, by increasing bond dimensions $D=400\to 1600$, the precision again gets continuously improved from $10^{-2}$ to $10^{-4}$.
@@
In Fig. 2 (c), we show the relative errors of the specific heat with respect to the ED results. It can be seen that, at low-temperature regime $T\lesssim 10^{-1}$, the 2S and CBE results exhibit similar accuracies for various bond dimensions from 400, 800, to 1600, which proves its validity in the considered projected Hamiltonian. In Fig. 2 (d), we compare the computational complexity of these two types of simulations (2S and CBE). Overall, the CBE simulations consume much less CPU time compared to the 2S simulations. At small bond dimensions, due to the numerical overheads, the acceleration of CBE scheme is below the ideal factor $d^{2}=4$. But as the bond dimensions increase, the speed up effect of CBE is getting increasingly pronounced; i.e., at the largest bond dimension $D=1600$ consindered, the acceleration is over a factor of $3$. Overall, we find tanTRG combined with CBE update scheme provides an efficient and high-precision approach to the study of the finite-temperature properties of the LLL-projected Hamiltonians.
@@
Finite-temperature properties of Composite Fermi Liquid.— We then move on to large system sizes with various number of orbitals $N=24,36,48$ with a fixed circumference $L_{y}=12$ and a relatively long screening length $\lambda=10$. As shown in Fig. 3 (a), we calculate the specific heat $c_{\mathrm{V}}$ down to low temperatures $T\lesssim 10^{-2}$. For the following analysis, we exclude the lower-$T$ regime data as depicted by the grey area. This is suggested by the compressibility $\partial\langle n\rangle/\partial\mu$ data [c.f. SM [83]] which quickly vanishes at $\sim 10^{-2}$, indicating a finite gap due to the finite circumference. Above the grey regime, the three specific heat curves exhibit power-law behaviour $T^{\alpha}$, and all of the three curves apparently deviate from the Fermi liquid behaviour of linear-$T$ scaling. As shown in Fig. 3 (b), notably, for the two larger sizes $N=36$ and $48$, the power $\alpha$ converge to the value of $2/3$, which is consistent with the value of $2/3$ predicted by Halperin, Lee, and Read (HLR) [42]. In the plot, we have also shown the smaller $\lambda=2$ data, they seem to suffer severer finite-size effect, while the power still tend to approach to $2/3$. The fact that such a non-Fermi-liquid thermodynamic response – different from the $C\sim T$ of normal metal, is the decisive evidence that there exists strong interactions among the CFs, and such interactions are mediated by the emergent gauge field in the QED3 description [42, 47, 48, 49, 50].

### v2

@@ Abstract
The physics of two-dimensional electron gas in a perpendicular magnetic field, i.e., the quantum Hall system, is remarkably rich. At half filling of the lowest Landau level, it has been predicted that “composite fermions”—emergent quasiparticle of an electron with two magnetic flux quanta—experience zero net magnetic field and form a Fermi sea, dubbed composite Fermi liquid (CFL). However, the seemingly simple appearance of CFL is a strongly correlated quantum many-body state in disguise, and to solve it in a controlled manner is extremely difficult, to the level that the thermodynamic properties of CFL is still largely unknown. In this work, we perform state-of-the-art thermal tensor network simulations on the $\nu=1/2$ Landau level systems, and observe low-temperature power-law behaviour of the specific heat, signaling the gapless nature of CFL. More importantly, the power is extracted to be closed to $2/3$, clearly deviated from the ordinary linear-$T$ Fermi liquid behaviour, suggesting the coupling between the CFs and the dynamical emergent gauge field and therefore revealed the quantum many-body aspect of the CFL state. Relevance of our methodology to other quantum Hall settings and moiré systems is discussed.
@@
Introduction.— The two-dimensional electron gas subjected to a perpendicular magnetic field—known as the quantum Hall (QH) system—exhibits remarkable physical richness. In the integer quantum Hall regime, the Hall conductance $\sigma_{xy}$ is quantized precisely in integer multiples of $e^{2}/h$ [1]. In the fractional quantum Hall regime, $\sigma_{xy}$ quantizes at fractional values of $e^{2}/h$, accompanied by the emergence of fractionally charged quasiparticles obeying abelian fractional statistics [2, 3, 4]. Notably, at filling $\nu=1/2$, the system exhibits metallic behavior, characterized by a finite longitudinal resistivity $\rho_{xx}$ and the absence of $\rho_{xy}$ pleateau  [5].
@@
The composite fermion (CF) theory postulates that electrons attach an even number of magnetic flux quanta, forming emergent quasiparticles that experience a reduced effective magnetic field [6, 7]. Along the line of this flux-attachment concept, mean-field theories [8, 9, 10, 11] are successful: near $\nu=1/2$, ballistic transport experiments are consistent with composite fermions moving in a weak residual magnetic field [12, 13, 14]; at half-filling, the external magnetic field is entirely canceled by the attached flux, resulting in a Fermi sea of composite fermions—a picture corroborated by surface acoustic wave measurements [15]. More recently, this conventional CF picture has been revisited in light of connections to 3D topological insulators (TIs), reformulating composite fermions as massless Dirac fermions reminiscent of TI surface states [16, 17, 18, 19]. This “composite Dirac fermion” theory couples fermions to an emergent gauge field, exhibiting non-Fermi-liquid behavior and providing a field-theoretic realization of (2+1)-dimensional quantum electrodynamics (QED3) as a strongly coupled conformal field theory, with both condensed matter [20, 21, 22, 23, 24, 25, 26, 27, 28] and high-energy [29, 30, 31, 32, 33, 34, 35, 36] physics correspondence.
In recent years, the fractional Chern insulator (FCI) [37, 38, 39, 40, 41, 42]—a lattice analogue of the fractional quantum Hall effect (when there is no mangetic field the state is also denoted as fractional quantum anomalous Hall (FQAH) effect)—has attracted considerable interest in the context of two-dimensional moiré materials, with extensive theoretical explorations [43, 44, 45, 46, 47, 48, 49, 50, 51]. FCIs are reported at high magnetic field in early experiment [52] at the fractional filling of Hofstadter band [53], and the fractional filling of the native Chern band in magic-angle twisted bilayer graphene (TBG) with the finite magnetic field to improve the quantum geometry, i.e., smoothing the Berry-curvature distribution and reducing violations of the trace condition of the ideal Landau-level-like relation $\operatorname{Tr}(g(\mathbf{k}))=|\Omega(\mathbf{k})|$ [54, 55]. More recently, the zero-field FCIs, i.e. FQAHs, are observed in twisted bilayer MoTe2 (tMoTe2) and rhombohedral graphene/hBN superlattices [56, 57, 58, 59, 60]. Notably, the zero-field metallic state at $\nu=1/2$ of the topological bands have also been theoretically and experimentally explored [61, 62, 57, 60].
@@
Substantial numerical research are dedicated to understanding the gapped FQH as well as the zero-field FCI phases, including the time-dependent variational principle [63, 64, 65] and single mode approximation [66] concerning the spectral properties, and thermal tensor network simulations [67] on the thermodynamical properties. In contrast, the numerical characterization of the half-filled CFL remains challenging. From this perspective, representative works include density matrix renormalization group (DMRG) investigation on the role of particle-hole symmetry  [68], variational Monte Carlo studies on the scaling of entanglement entropy [69, 70, 71], effects of anisotropy [72], DMRG studies on the relationship between the Fermi sea of bare electrons at zero field and that of composite fermions at high field [73] as well as the influence of discrete rotational symmetry $C_{N}$ [74], and exact diagonalization calculation on the thermoelectric response [75]. In all these directions, new computation and discussions of often-time contradicting results are actively going on.
@@
In this work, we investigate the thermodynamic properties of the composite Fermi liquid at half filling of the lowest Landau level using large-scale thermal tensor network simulations, specifically the tangent-space tensor renormalization group (tanTRG) method [76]. We observe a low-temperature specific heat that scales as $T^{\alpha}$, with $\alpha\approx 2/3$—indicative of gapless excitations and markedly inconsistent with the linear-$T$ dependence expected for a conventional Fermi liquid. This power-law behavior provides direct evidence for the strongly correlation nature of CFL state, in that, the coupling between composite fermions and a dynamical emergent gauge field gives rise to non-Fermi liquid properties. In a broader context, our results put forward the theoretical framework that such a non-Fermi-liquid could emerge from Fermi surface coupled to a critical ($U(1)$) gauge field [25, 26], ferromagnetic bosonic fluctuations [77, 78] and nematic fluctuations [79]. And our thermal tensor-network method can be straightforwardly generalized to FQAH and moiré systems, where the continuum Hamiltonian is naturally defined in momentum space.
@@
where $V(\mathbf{r}_{1},\mathbf{r}_{2})\equiv\frac{e^{-|\mathbf{r}_{1}-\mathbf{r}_{2}|/\lambda}}{|\mathbf{r}_{1}-\mathbf{r}_{2}|}$ is the Coulomb-Yukawa interaction between electrons, $n(\mathbf{r})\equiv c^{\dagger}(\mathbf{r})c(\mathbf{r})$ is the particle number operator and the factor $\tfrac{1}{2}$ is utilized to keep the half-filling condition in the grand-canonical ensemble. We then project the Hamiltonian onto the lowest Landau level (LLL) via $c(\mathbf{r})=\sum_{n}\phi_{n}(\mathbf{r})c_{n}$, where $\phi_{n}(\mathbf{r})\propto e^{-\frac{1}{2}(\frac{x-x_{n}}{\ell})^{2}}e^{\mathrm{i}k_{n}y}$ is the LLL wavefunction in the cylindrical geometry (i.e., Landau gauge) with $k_{n}=2\pi n/L_{y}$, $x_{n}=2\pi n\ell^{2}/L_{y}$, $n\in[0,N-1]$ and $N$ the number of LLL orbitals [c.f. Fig. 1 (a)]. We set the magnetic length $\ell=\sqrt{\hbar c/(eB)}=1$ and $k_{\mathrm{B}}=1$ as the unit throughout. This yields the LLL-projected Hamiltonian,
@@
with the form factor $\mathcal{A}_{m,n,k,l}=\int\mathrm{d}r_{1}\mathrm{d}r_{2}V(r_{1}-r_{2})\phi^{\ast}_{m}(r_{1})\phi_{l}(r_{1})\phi^{\ast}_{n}(r_{2})\phi_{k}(r_{2})$ being evaluated numerically.
Tangent-space tensor renormalization group (tanTRG) in the LLL basis.— Here, we briefly recapitulate the idea of tangent-space tensor renormalization group (tanTRG) which is proved to be successful in accessing the challenging low-temperature properties of quantum lattice models, e.g., 2D Hubbard model and $t$-$J$ model for cuprate [76, 80], bilayer $t$-$J$-$J_{\perp}$ model for Nickelate superconductor [81], and real-space effective model of MoTe2 [82]. We then show how tanTRG can be adapted to the LLL orbital basis.
To begin with, a generic Hamiltonian can be efficiently decomposed into matrix product operator (MPO) representation [83, 84, 85, 86, 87]. At high temperature, the themal density operator can then be constructed via series expansion with high precision [88], which effectively reduce to $\rho(\tau)\equiv e^{-\tau\mathcal{H}}=I-\tau\mathcal{H}$ for sufficiently small inverse temperature $\tau\equiv 1/T$ (here $\tau=10^{-6}$ is taken rendering a small expansion error of $\mathcal{O}(10^{-12})$).
We then cool down the system through a mixed temperature grid, i.e., firstly a exponential one $\beta_{i}=\tau\times 2^{i/z}$ to quickly approaching the low-$T$ regime (here $z=4$ to ensure dense temperature point for numerical derivative), which then follows by a linear (inverse) temperature grid when the inverse temperature increase $\Delta\beta$ exceeds a given threshold (here $\Delta\beta=5$ is taken) to ensure small projection errors in low-$T$ regime.
The cooling step is followed by the flow equation $\partial\rho/\partial\beta=\mathcal{P}(-\mathcal{H}\rho)\mathcal{P}$ constraint to a fixed-bond-dimension matrix product operator (MPO) manifold $\mathcal{M}_{\mathrm{MPO}}$, with $\mathcal{H}\rho$ being the tangent vector in the full Hilbert space which generally takes the density matrix $\rho$ out of $\mathcal{M}_{\mathrm{MPO}}$, and the projector $\mathcal{P}$ serves to projecting the tangent vector back to the tangent space of the MPO manifold $\mathcal{T}_{\rho}\mathcal{M}_{\mathrm{MPO}}$. For more details concerning the MPO parameterization of the above constraint flow equation, we refer the interested readers to Ref. [76].
@@
Specifically, in this work, we deal with the many-body basis consisted of LLL wavefunctions, $|q_{0},q_{1},\cdots,q_{N-1}\rangle=(c^{\dagger}_{0})^{q_{0}}(c^{\dagger}_{1})^{q_{1}}\cdots(c^{\dagger}_{N-1})^{q_{N-1}}|\Omega\rangle$, with the quantum number $q_{i}$ denoting the number of particles filled in the $i$th LLL orbital, and $|\Omega\rangle$ being the vacuum state. A generic operator $O$ in this many-body LLL basis then express as $O=\sum_{\{q_{i},q_{i}^{\prime}\}}O_{\{q_{i}\},\{q_{i}^{\prime}\}}|\{q_{i}\}\rangle\langle\{q_{i}^{\prime}\}|$, where $O_{\{q_{i}\},\{q_{i}^{\prime}\}}\equiv\langle\{q_{i}\}|O|\{q_{i}^{\prime}\}\rangle$ is a $2^{N}\times 2^{N}$-dimensional tensor, and can be parameterized into the MPO form, $O_{\{q_{i}\},\{q_{i}^{\prime}\}}=\sum_{b_{0},b_{1},\cdots,b_{N}}\prod_{i=0}^{N-1}A_{b_{i},b_{i+1},q_{i},q_{i}^{\prime}}$ with the virtual indices $b_{i}$’s ranging from $1$ to $D$ (bond dimension), as shown in Fig. 1 (b). For the projected Hamiltonian $\mathcal{H}^{\mathrm{LLL}}$ and the initial high-temperature density operator $\rho(\tau)$, we can find the exact MPO forms for them, and thus the lower-temperature density operator $\rho(\beta)$ can then be tracked with the tanTRG protocol as mentioned above. We note that our method can be straightforwardly generalized to FQAH and quantum morié systems [89, 43, 44, 45, 46, 47, 48, 49, 50, 51, 90] and other projected Hamiltonian [91, 92] systems, where one only needs to change the basis function on which the projection is carried out.
@@
Benchmark results.— In Fig. 2, we benchmark tanTRG in a small system size $N=16$. Here, we take the circumference of the cylinder $L_{y}=12$, and the effective screening length $\lambda=10\tfrac{2\pi\ell^{2}}{L_{y}}$. We firstly calculate the free energy per orbital $f(T)\equiv-\frac{1}{N}T\ln Z$ with $Z=\operatorname{Tr}\rho$ the partition function.
In Fig. 2 (a), we have shown three curves of the free energy relative errors $\delta f\equiv|f_{\mathrm{tanTRG}}-f_{\mathrm{ED}}|/|f_{\mathrm{ED}}|$ with incremental bond dimensions $D=400,800,1600$. For all the three curves, at the high-temperature regime, the relative errors remain within the order of $10^{-6}$ as the discarded weights can be neglected and the final error comes from the Trotter error and the projection error (the one results from projecting the full tangent vector into $T_{\rho}\mathcal{M}_{\mathrm{MPO}}$). At intermediate temperature $T\sim 10^{-1}$, the thermal state changes significantly as indicated by the round peak of the specific heat [c.f. the inset of Fig. 2 (b)], and the finite discarded weights set in, resulting in a relatively higher errors of free energies. At lower-temperature regime, finite-size gap sets in and the thermal density operator $\rho$ converges to $|\psi_{0}\rangle\langle\psi_{0}|$ (ground state) which in this example possess lower (purified) entanglement entropy thus lower $\delta f$ relative to the intermediate-$T$ cases. Such lower-$T$ behaviour can also be understood from the MPO manifold perspective, which becomes more flattened when the finite-size gap sets in and renders small projection errors. Furthermore, by increasing $D$, the precision increases from $10^{-3}$ to $10^{-5}$ accordingly in the low-$T$ regime.
In Fig. 2 (b), we present the relative error of specific heat $c_{V}\equiv\partial e/\partial T$, with $e\equiv\frac{1}{N}\operatorname{Tr}(\mathcal{H}\rho)/Z$ being the internal energy per orbital. Compared with the exact diagonalization results, the relative errors of specific heat $\delta c_{V}$ is shown as function of temperature $T$. Similar to the free energy, the specific heat $c_{V}$ possesses high precision with relative errors within $10^{-8}$ for high temperatures $T\gtrsim 1$, which increase quickly at $T\sim 10^{-1}$, and at lower temperatures $\delta c_{V}$ even decreases slightly due to existence of the finite-size gap. Furthermore, by increasing bond dimensions $D=400\to 1600$, the precision again gets improved from $10^{-2}$ to $10^{-4}$.
@@
In Fig. 2 (c), we show the relative errors of the specific heat with respect to the ED results. It can be seen that, at low-temperature regime $T\lesssim 10^{-1}$, the 2S and CBE results exhibit similar accuracies for various bond dimensions from 400, 800, to 1600, which proves its validity in the considered projected Hamiltonian. In Fig. 2 (d), we compare the computational complexity of these two types of simulations (2S and CBE). Overall, the CBE simulations consume much less CPU time compared to the 2S ones. At small bond dimensions, due to the numerical overheads, the acceleration of CBE scheme is below the ideal factor $d^{2}=4$. But as the bond dimensions increase, the speed up effect of CBE is getting increasingly pronounced; i.e., at the largest bond dimension $D=1600$ consindered, the acceleration is over a factor of $3$. Overall, we find tanTRG combined with CBE update provides an efficient and high-precision approach to the study of the finite-temperature properties of the LLL-projected Hamiltonians.
@@
Finite-temperature properties of Composite Fermi Liquid.— We then move forward to larger system sizes with various number of orbitals $N=24,36,48$ with a fixed circumference $L_{y}=12$ and a relatively long screening length $\lambda=10$. As shown in Fig. 3 (a), we calculate the specific heat $c_{V}$ down to low temperatures $T\lesssim 10^{-2}$. For the following analysis, we exclude the lower-$T$ regime data as depicted by the grey area to minimize the finite-size effect. This is suggested by the compressibility $\partial\langle n\rangle/\partial\mu$ data [c.f. SM [95]] which quickly vanishes at $\sim 10^{-2}$, indicating a finite-circumference gap. Above the grey regime, the three specific heat curves exhibit power-law behaviour $T^{\alpha}$, and they all apparently deviate from the linear-$T$ Fermi-liquid behaviour. As shown in Fig. 3 (b), notably, for the two larger sizes $N=36$ and $48$, the power $\alpha$ converge to the value of $2/3$, consistent with the value predicted by Halperin, Lee, and Read (HLR) [11]. In the plot, we have also shown the smaller $\lambda=2$ data, they seem to suffer severer finite-size effect, while the power still tend to approach to $2/3$. The fact that such a non-Fermi-liquid thermodynamic response – different from the $C\sim T$ of normal metal, is the decisive evidence that there exists strong interactions among the CFs, and such interactions are mediated by the emergent gauge field in the QED3 description [11, 16, 17, 18, 19, 25, 26], or ferromagnetic and nematic critical bosonsic fluctuations [79, 77, 78].